OpenStudy (anonymous):

Solve. 8 over x + 15 over x^2 = -1

OpenStudy (anonymous):

Multiply both sides by x^2. You will get a quadratic then :-)

OpenStudy (anonymous):

i dont know how can you help

OpenStudy (anonymous):

You have \( \frac{8}{x} + \frac{15}{x^2} = -1 \) multiply both sides by \( x^2\), you should get: \[ x^2 (\frac{8}{x} + \frac{15}{x^2}) = -x^2\]Can you simplify it?

OpenStudy (anonymous):

im sorry i really dont understad

OpenStudy (anonymous):

You don't understand multiplying both sides by the same number? Something like 2 + 3 = 5 is equal to 2(2+3) = 10?